# Is Grokking a Computational Glass Relaxation?

### v1

@@ 6 Conclusions and Discussion
What causes the NN to fall into the memorization state? One requirement for the formation of a glass is to cool the liquid below the glass transition temperature hansen2013theory . Below this temperature, the free movement of atoms is greatly restricted, thus hindering the system from evolving towards equilibrium. In grokking tasks such as modular arithmetic tasks, the NNs are severely overparameterized, i.e., the number of the parameters is much larger than the size of the dataset. During training, the NN can fully memorize all training data and achieve a very low training loss. If we take the training loss as the energy, this is equivalent to lowering the temperature below the glass transition temperature. Another requirement for glass formation is a sufficiently fast cooling rate, so that the particles do not have time to rearrange themselves into the high-entropy crystalline structure. During the training process, the gradient descent based optimizer, especially adaptive optimizer like Adam, directs the parameters of the NN to move in the direction of decreasing training loss with only very small perturbations (caused by large learning rate, small batch size, and data accuracy, etc.), making it easier to form computational glass. In contrast, the Langevin dynamics in the WanD optimizer we designed enriches the perturbations in training, thus avoiding the formation of computational glass.
What is the internal dynamics of the grokking process? The non-equilibrium glass formed by rapidly cooling the liquid can move toward equilibrium through a kinetic process, which is called structural relaxation riechers2022predicting . Since atoms move very slowly at low temperatures, the relaxation time can be very long relative to the cooling process. Our results show that in grokking tasks, the equilibrium state at low training loss generalizes well. For adaptive-learning-rate optimizers such as Adam, the parameters in the basin of the loss landscape have very small gradients, so there is not enough speed to update the NN parameters towards the equilibrium state, which causes a significant time delay from memorization to generalization. Although the time of the grokking process varies in multiple random trainings runs, the generalization process is always much longer than the memorization process power2022grokking , similar to the relaxation of glasses riechers2022predicting .
@@
Our conclusions agree and disagree with some previous studies. There are a series of work argued that grokking is due to a transition from poor to good representation learning kumar2023grokking ; varma2023explaining ; kunin2024get . This agrees with our analogy of relaxation from a poorly structured “glassy” (memorizing) state to a more ordered (generalizing) configuration. In fact, our entropy perspective offers a resolution to debates regarding solution “efficiency” and parameter norms in previous works kumar2023grokking ; varma2023explaining . While Varma et al.varma2023explaining argue that grokking requires a transition to an “efficient” generalizing circuit characterized by a lower parameter norm, Kumar et al.kumar2023grokking challenges it as they identified generalizing solutions with higher norm than memorizing solutions. Our framework suggests this conflict stems from using parameter norm alone to define efficiency and entropy is a better indicator since the efficient solution will have more “free” parameters and therefore a larger entropy. This is also consistent with Merrill et al. merrill2023tale who showed that memorization solution and generalization solution correspond to dense and sparse subnetworks, respectively. In the sparse subnetworks, only a few neurons participate in the memorization of the rules, and the remaining degrees of freedom contribute to the entropy, so it should have a larger entropy as we observed. Kumar et al. kumar2023grokking proposed that the key determinant of grokking is the rate of feature learning, which is consistent with our glass analogy, where the rate-limited free motion of particles is an important reason for the glass formation and long relaxation time. Our work also supports the finding that perturbation is beneficial to the grokking learning process thilak2022slingshot . Rubin et al. rubin2023grokking claimed that grokking is a first-order phase transition process, but our results overturn this conclusion. This is probably because their network architecture and training data are different from ours, while ours are consistent with the original paper that defines grokking. In addition, the high-weight norm generalization states we found through WanD also challenge the conclusions of Liu et al. liu2022omnigrok .
@@
Our hypothesis of grokking and glass relaxation invites further investigation. While we focused on experiments with the original grokking tasks power2022grokking , future studies should examine whether our findings are applicable for other grokking tasks. Furthermore, a recent study has compared the behavior of deep neural network and structural glass with quantitative measurements winter2025glassy , and we can refer to its research method to verify whether the grokking NN is a structural glass. In addition, it remains unclear how does glassy dynamics affect the effectiveness of regularization methods in removing grokking. Finally, we also aim to further improve the WanD optimizer to enhance its stability and general applicability. Techniques from statistical physics designed to accelerate equilibration, such as simulated annealing and parallel tempering, can also be applied to WanD’s future refinement frenkel2023understanding .
@@ A.2 Training Details and Cost
For modular arithmetic tasks training by AdamW, we use a full batch size on 1 GeForce RTX 4090 GPU, training for 10000 optimization steps with weight decay 1 or fixed weight norm 30. The hyperparameters for training with AdamW are given in Table 1. For our toy optimizer WanD training, since it is not yet fully developed, there are many hyperparameters that directly affect the generalization of the training results. We provide a set of hyperparameter combinations in the code that can generalized well.
@@ Appendix B Additional Results
[…]
@@
[…]

### v2

@@ 6.2 Discussion
What causes the NN to fall into the memorization state? One requirement for the formation of a glass is to cool the liquid below the glass transition temperature hansen2013theory . Below this temperature, the free movement of atoms is greatly restricted, thus hindering the system from evolving towards equilibrium. In grokking tasks such as modular arithmetics, the NNs are severely overparameterized, i.e., the number of the parameters is much larger than the size of the dataset. During training, the NN can fully memorize all training data and achieve a very low training loss. If we take the training loss as the energy, this is equivalent to lowering the temperature below the glass transition temperature. Another requirement for glass formation is a sufficiently fast cooling rate, so that the molecules do not have time to rearrange themselves into the lower-free-energy crystalline structure. During the training process, the gradient descent based optimizers, especially adaptive optimizer like Adam, effectively implement a rapid ’cooling’ process by aggressively driving parameters towards regions of reduced training loss. This rapid descent often proceeds with insufficient perturbations (caused by large learning rate, small batch size, and data accuracy, etc.) to escape from poor local minima, making NNs easier to reside in a non-generalizable loss minima and therefore form a computational glass. In contrast, the Langevin dynamics in the WanD optimizer we designed enriches the perturbations in training, thus avoiding the formation of computational glass.
What is the internal dynamics of the grokking process? When a liquid is rapidly cooled, it can form a non-equilibrium glassy state that gradually approaches equilibrium through structural relaxation—a slow kinetic process riechers2022predicting . Due to the sluggish motion of atoms at low temperatures, the relaxation time can far exceed the timescale of the cooling process. Analogously, in grokking tasks, the equilibrium state at low training loss corresponds to good generalization. However, parameters located in the low train loss regime exhibit very small gradients, which severely limits the ability of gradient-based optimizers to effectively update the network toward this generalizing equilibrium. This dynamical slowdown results in a substantial delay between the memorization and generalization phases. The effect is further exacerbated by adaptive optimizers like Adam, where the non-uniform scaling of gradients can further dampen parameter updates, increasing the likelihood of being trapped in non-generalizable minima zhou2020towards . Although the duration of the grokking process varies across training runs, as observed in prior studies and our experiments, the transition to generalization consistently takes significantly longer than the memorization stage power2022grokking , mirroring the slow relaxation behavior characteristic of glassy systems riechers2022predicting .
@@
Our conclusions agree and disagree with some previous studies. There were a series of work arguing that grokking is due to a transition from poor to good representation learning kumar2023grokking ; varma2023explaining ; kunin2024get ; liu2022towards . This agrees with our analogy of relaxation from a poorly structured “glassy” (memorizing) state to a more ordered (generalizing) configuration. In fact, our entropy perspective offers a resolution to debates regarding solution “efficiency” and parameter norms in previous works kumar2023grokking ; varma2023explaining . While Varma et al.varma2023explaining argue that grokking requires a transition to an “efficient” generalizing circuit characterized by a lower parameter norm, Kumar et al.kumar2023grokking challenges it as they identified generalizing solutions with higher norm than memorizing solutions. Our framework suggests this conflict stems from using parameter norm alone to define efficiency, while entropy is actually a better indicator since the efficient solution will have more “free” parameters and therefore a larger entropy. This is also consistent with Merrill et al. merrill2023tale who showed that memorization solution and generalization solution correspond to dense and sparse subnetworks, respectively. In the sparse subnetworks, only a few neurons participate in the memorization of the rules, and the remaining degrees of freedom contribute to the entropy, so it should have a larger entropy as we observed. Kumar et al. kumar2023grokking proposed that the key determinant of grokking is the rate of feature learning, which is consistent with our glass analogy, where the rate-limited free motion of particles is an important reason for the glass formation and long relaxation time. The finding by Liu et al. liu2022towards that a fast-learning decoder can help trigger grokking also supports our analogy, as the rapid decoder acts similarly to ‘fast cooling’, trapping the network in a glassy memorization state. Our work also supports the finding that perturbation is beneficial to the grokking learning process thilak2022slingshot , as WanD is able to eliminate grokking directly. Rubin et al. rubin2023grokking claimed that grokking is a first-order phase transition process, but our results overturn this conclusion. This is probably because their network architecture and training data are different from ours, while ours are consistent with the original paper that discovered grokking. In addition, the high-weight norm generalization states we found through WanD also challenge the conclusions of Liu et al. liu2022omnigrok . Due to space limitations, we provide additional discussion in the Appendix C.
6.3 Limitation and future work
Limitation: i). Task scope: Our study is limited to one-layer transformers on modular arithmetic tasks and does not include other grokking benchmarks (e.g., tasks in rubin2023grokking ; liu2022omnigrok ). ii). Interaction with explicit regularization: The relationship between “glassy dynamics” and various explicit regularization methods in grokking systems remains unclear and requires further exploration. iii). Proof-of-concept nature of WanD optimizer: WanD has only been tested in arithmetic task scenarios, is sensitive to hyperparameters, and exhibits instability near generalization.
@@ 6.3 Limitation and future work
Our hypothesis of grokking and glass relaxation invites further investigation. While we focused on experiments with the original grokking tasks power2022grokking , future studies should examine whether our findings are applicable for other grokking tasks. Furthermore, a recent study has compared the behavior of deep neural network and structural glass with quantitative measurements winter2025glassy , and we can refer to its research method to verify whether the grokking NN is a structural glass. In addition, it remains unclear how does glassy dynamics affect the effectiveness of regularization methods in removing grokking. Finally, we also aim to further improve the WanD optimizer to enhance its stability and general applicability. As an algorithm based on WLMD, WanD involves multiple hyperparameters that require tuning. Our experiments identify the learning rate, temperature, and friction coefficient as the three most influential factors affecting training outcomes. We plan to further investigate empirical guidelines for effective hyperparameter selection. Techniques from statistical physics designed to accelerate equilibration, such as simulated annealing and parallel tempering, can also be applied to WanD’s future refinement frenkel2023understanding . We anticipate that approaches such as a dynamic hybrid optimization strategy combining WanD and AdamW could help mitigate instability issues and present some preliminary results in the Appendix B.3. Ultimately, we aim to develop WanD into a general-purpose optimizer applicable to a broader range of machine learning tasks.
Acknowledgments
The authors thank National Natural Science Foundation of China for supporting this research (Grant 12405043, G.Z.).We also thank computational resources provided by Bridges-2 at Pittsburgh Supercomputing Center through ACCESS allocation CIS230096 (E.Y.).
@@ A.2 Training Details and Cost
For modular arithmetic tasks training by AdamW, we use a full batch size on 1 GeForce RTX 4090 GPU, training for 10000 optimization steps with weight decay 1 or fixed weight norm 30. The hyperparameters for training with AdamW are given in Table 1. For our toy optimizer WanD training, since it is not yet fully developed, there are many hyperparameters that directly affect the generalization of the training results. Among them, the learning rate, temperature, and friction coefficient are the three key factors that determine the training results. We provide a set of hyperparameter combinations in the code that can generalized well.
@@
A.3 WLMD details
Smoothed accuracy: To ensure well-defined gradient of test accuracy with respect to NN parameters in WLMD, we compute a smoothed test accuracy using a sigmoid function, following [2]. The detailed procedure for obtaining smoothed accuracy is outlined in Algorithm 2. In our work, we choose the smoothed accuracy slope $s=7$.
Algorithm 2 Smoothed Accuracy
Transformer model $T(\cdot)$, dataset $(X=\{x_{1},x_{2},...,x_{n}\},\hat{Y}=\{\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{n}\})$, smoothed accuracy slope $s$
The value of smoothed accuracy close to the true accuracy, and the gradient of smoothed accuracy with respect to parameters has a finite value.
[…]
  Get the highest logit among all wrong output classes: $p_{i}(y=y_{i\_wrong})$ where $y_{i\_wrong}\in Y_{i}=T(x_{i})$
  Get the logit of the correct output class: $p_{i}(y=\hat{y}_{i})$
[…]
end for
[…]
Entropy bias: When sampling the entropy landscape of grokking NNs, we are interested in the regime with lower training loss, which corresponds to the memorization state and the generalization state. However, areas with high training loss usually have higher entropy (as evidenced by randomly initialized networks that generally possess high training loss), which will waste a lot of time exploring in the early stages of sampling. In order to prevent the simulation program from wasting time outside the range of our interest, we set an additional entropy bias:
$\mathrm{if}\ ln(L_{train})>L_{0},\ S_{bias}(ln(L_{train}))=\beta(ln(L_{train})-L_{0})^{2};\ \mathrm{else}\ S_{bias}(ln(L_{train}))=0.$ (7)
If the training loss is outside the range of interest $(ln(L_{train})>L_{0})$ when updating the NN parameters, this entropy bias will provide additional entropy gradients to to force the NN to move towards lower training loss.
The overall WLMD entropy sampling procedure is outlined in Algorithm 3. We employ the cross-entropy loss for our modular arithmetic task and a logarithmic scale of the training loss is used to better examine the entropy landscape near zero training loss.
Algorithm 3 WLMD Entropy Sampling Method
[…]
Convergent entropy landscape $S(x_{WL},y_{WL})$ where $x_{WL}=ln(L_{train})$, $y_{WL}=A_{test}$;
Initialize: $S\leftarrow S_{bias}$;
for each sampling epoch do
  Compute the logarithm of the train loss and smoothed test accuracy:
  $ln(L_{train})=ln(CrossEntropyLoss(T(X_{train}),y_{train}))$;
  $A_{test}=SmoothedAccuracy(T(X_{test}),y_{test})$;
  Compute entropy gradient $x_{grad}=\frac{\partial S(x_{WL},y_{WL})}{\partial x_{WL}}$ and $y_{grad}=\frac{\partial S(x_{WL},y_{WL})}{\partial y_{WL}}$;
  The combined loss $L_{combined}=-x_{WL}x_{grad}-y_{WL}y_{grad}$;
  Do $L_{combined}.backward()$ and get $\frac{d^{2}c_{i}}{dt^{2}}$;
  for each parameter $c_{i}$ do
   Update the parameter velocity by Equation 5: $\frac{dc_{i}}{dt}\leftarrow\frac{dc_{i}}{dt}(1-c_{f})+\frac{d^{2}c_{i}}{dt^{2}}\Delta t+v_{Langevin}$;
   Update the parameter: $c_{i}\leftarrow c_{i}+\frac{dc_{i}}{dt}\Delta t$;
  end for
  Update the Wang-Landau factor by Equation 3: $f_{WL}\leftarrow F_{WL}min\{t/t_{WL},t_{WL}/t\}$;
  Update the entropy landscape by Equation 3: $S(x_{WL},y_{WL})\leftarrow S(x_{WL},y_{WL})+f_{WL}\delta(x-x_{WL},y-y_{WL})$;
end for
$S\leftarrow S-S_{bias}$;
@@ B.1 Visualization of modular arithmetic tasks
Figure 4: Visualization of three different arithmetic tasks. (a) x+ymod 67x+y\ mod\ 67. (b) $x^{2}+y\ mod\ 67$. (c) $x^{3}+xy^{2}+y\ mod\ 67$. Our protocol is equivalent to randomly taking 50% of the pixels from the image as the training set and trying to complete the rest. The more complex the expression, the more visually irregular the image is.
B.2 AdamW training results of modular arithmetic tasks
@@
Figure 5: Training results of three different arithmetic tasks: (a) x+ymod 67x+y\ mod\ 67 is easy to learn, but the hysteresis of the test accuracy shows grokking. (b) $x^{2}+y\ mod\ 67$ is hard to learn, the grokking is more obvious. (c) $x^{3}+xy^{2}+y\ mod\ 67$ cannot be learned.
@@
B.3 Training results of AdamW and WanD under different dataset splits
Increasing the training dataset split fraction is an effective way to eliminate grokking [8]. We tested the training results of AdamW and WanD on the x+ymod 67x+y\ mod\ 67 (Figure 7) and $x^{2}+y\ mod\ 67$ (Figure 8) tasks at the same learning rate 0.003, with the training dataset fraction being 50%, 60%, 70%, and 80%, respectively. Here we anneal WanD with AdamW so that the update weight of WanD increases first and then decreases during training, thus avoiding the instability problem at the end of WanD training. Specifically, in each iteration, the NN parameters are updated by the WanD optimizer with weight $\alpha$ and by the AdamW optimizer with weight $1-\alpha$, where $\alpha=(2A_{test}-1)^{2}$.
Figure 7: AdamW and WanD training results of x+ymod 67x+y\ mod\ 67 task with different training dataset fraction. Top row:AdamW; bottom row:WanD; (a)(e) 50%, (b)(f) 60%, (c)(g) 70%, and (d)(h) 80% training dataset split fraction. The hyperparameter of WanD: kT=0.5, frictionCoefficient=0.5, maxTrainLossToStudy=-8.
Figure 8: AdamW and WanD training results of $x^{2}+y\ mod\ 67$ task with different training dataset fraction. (a)(b)(c)(d):AdamW; (e)(f)(g)(h):WanD; (a)(e)/(b)(f)/(c)(g)/(d)(h): 50%/60%/70%/80% training dataset split fraction. The hyperparameter of WanD: kT=0.7, frictionCoefficient=0.6, maxTrainLossToStudy=-8. It is worth noting that WanD in (e) achieves generalization with nearly three times the efficiency of AdamW in (a).
The results show that WanD has a significant training efficiency advantage in tasks with severe grokking, and even in tasks where grokking is greatly eliminated, WanD can still achieve generalization ahead of AdamW. This result demonstrates the great potential of WanD as a toy optimizer and motivates us to continue improving it.
Appendix C Additional Discussion
Robustness of entropy landscape diagrams: In this work, we visualize the entropy landscape with respect to the logarithm of the training cross-entropy loss and the smoothed test accuracy. Due to the symmetry between training and test datasets, modifying the training loss formulation is conceptually equivalent to altering the test loss. Our reported results are based on the smoothed test accuracy, whose inverse can be interpreted as a form of loss function. We have also conducted internal experiments using the cross-entropy test loss in place of the smoothed accuracy. The resulting entropy diagram consistently support the same conclusion: entropy barrier is not observed, and only a single connected high-entropy region exists, indicating no sign of a first order phase transition. This consistency confirms that our findings are robust to the specific choice of the loss function.
Sparse subnetworks: As shown in [25], grokking is associated with sparse subnetworks in which only a subset of neurons is involved in rule memorization. From an entropy perspective, parameters outside these functional subnetworks do not affect generalization. Perturbing such parameters preserves model performance, implying that the corresponding solution set occupies a larger volume in parameter space. Since entropy is defined as the logarithm of this volume, neural networks with sparse subnetworks exhibit higher entropy.
Other entropy-based optimizer: WanD is an optimizer built on the one-dimensional WLMD framework. It excels at identifying high-entropy states, particularly the generalized solution, via molecular dynamics, leveraging the fact that generalized solutions possess higher entropy at low training loss. Prior to our work, there is another entropy-guided optimizers named Entropy-SGD [36] have been proposed. Below, we briefly highlight the theoretical distinctions between WanD and Entropy-SGD:
  • Different notions of entropy. The core distinction lies in the entropy definition: WanD targets the global Boltzmann entropy, reflecting the total volume of the solution space, whereas Entropy-SGD focuses on local entropy, which captures the flatness of individual minima.
  • Different methodologies. WanD estimates entropy via 1D WLMD sampling, while Entropy-SGD relies on Langevin dynamics to approximate local entropy gradients.
Optimizers like Adam often progress slowly due to limited gradient signals during the memorization-to-generalization transition [10]. In principle, Entropy-SGD could potentially accelerate grokking if its identified high local-entropy states exhibit strong generalization. A rigorous empirical comparison between these approaches remains a valuable direction for future work.